SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (13th Feb, 2007)

Attendees: Debbie Bard, Rainer Bartoldus, Al Eisner, Hojeong Kim, Selina Li, Rahmat Rahmat, Su Dong, Mark Tibbetts.

Operations

  • Paged Friday morning becasue trickle plots showed hole - updated long-term instructions to point out that this is NOT the trigger's domain.
  • Updated JAS pages to remove old DCT comparison plots, change contact names and highlight some key instructions.

Backgrounds

  • In the MDI meeting last week it was asked if the high IFR occupancies were trigger-related. It is extremely unlikely, but could be checked by looking at occupancies for runs where M*+5U was not the only trigger line.

EMT

  • Mark spoke to Jong - the problem mentionned was in comsics runs taken earlier this year where two unrelated areas in the EMC were showing systematically high occupancies. The way to test this is to look at pure DCH trigger events for those runs and see if the high occupancy persists.
  • Trackmatching study: the plot here shows efficiency vs cluster energy for track-matched clusters. This plot shows the phi co-ordinate of the cluster (x-axis) against the entry point of the track (y-axis).

GLT

  • The spare board that was found bad during the last test in IR2 has been fixed and looks good at the teststand. There was a problem with one, and possibly both dlink clocks.

L3

  • See the tracking group's hypernews discussion on the new tracking validation.
  • Rainer performed the new lumi calibration with the new bookeeping database with the new lumi numbers, and found almost exactly no change from the start-of-Run5 calibration. However, the step in lumi seen in August when the TDC bug was fixed has diaapoeared - has the sensitivity changed? Also, there's still the question of whether the hadron ration depends on luminosity. This could be looked into.

Data Quality and Monitoring

  • The Run 6 lumi stripchart (see p.119) shows a bit of a rise, but there are too few data points to be sure. Also, automatic y-axis scaling is not necceessarily to be desired.

Configuration

  • It seems to be impossible to get to the MC truth in the L1TNtupleMooseApp. Instead, Debbie could try making the collections with Moose and then running L1TKanApp over them. This will give GLT data at least. Or look for virtual compton events in data (L3 tag).
  • Joes is continuing to compare the Moose simulated trigger with the emulated trigger, but this isn't easy!

Who wants the trigger pager while Debbie's out of the country next week?

  • Well, no-one really, but Hojeong will take it and Selina will be on standby to help out with anything that comes up. Thankyou!

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.